OpenStudy (anonymous):

Help please ? A = (B + C)/D Solve for C. First, multiply both sides by D. A__ = B + C Then subtract both sides by B. A______ = C.

OpenStudy (anonymous):

@nirmalnema

OpenStudy (anonymous):

if you multiply both sides by D you get \(AD=B+C\)

OpenStudy (anonymous):

"subtract both sides by B" is not english, and therefore not math what it means is "subtract B FROM both sides" you get \(AD-B=C\)

OpenStudy (anonymous):

what grade are you in girl?

OpenStudy (anonymous):

Junior why ? & So thats my answer?

OpenStudy (nirmalnema):

A=B+C /D so AD = B+C therefore C = AD-B

OpenStudy (anonymous):

yes actually you only have to fill in the blanks right?

OpenStudy (anonymous):

yes

OpenStudy (nirmalnema):

YUP....

OpenStudy (anonymous):

put a D in the first blank do it looks like AD = B + C then the second line should look like AD - B = C

OpenStudy (anonymous):

Okie dokie thank you